OpenStudy (rachie19):

..please help... 2log(4)3x+log(4)5x=log(4)225

OpenStudy (campbell_st):

so they are all log base 4 so rewriting the equation using log laws you get \[\log_{4}(3x)^2 + \log_{4}(5x) = \log_{4}(225)\] addition in logs means multiplication \[\log_{4}(9x^2 \times 5x) = \log_{4}(225)\] which becomes \[\log_{4}(45x^3) = \log_{4}(225)\] since they are the same base you just need to solve \[45x^3 = 225\]

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

11 x Log[4] = 225 Log[4]